# Discrete-time thermodynamic speed limit

### v1

@@ V.3 System with unidirectional transition
$\mathbf{P}(t_{i+1})=[\mathbf{M}^{({\rm b})}+\mathbf{M}^{({\rm u})}]\mathbf{P}(t_{i}),$ (37)
@@
We evaluate time-reversed EP and dynamical activity for the one-time-step relaxation process described by Eq. (37). The results are plotted in Fig. 3. The dashed line represents the bound of the speed limit for a system including unidirectional transitions, as given by Eq. (33), using $h(x)=h_{\rm sym}(x)$. Data points indicate the results obtained from various values of $g$, $M_{1,0}$, and $M_{0,1}$ chosen within the range $(0.01,0.9)$ under the constraint $d_{\textrm{TV}}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{(u)}(t_{0})>0$. This example validates the speed limit for a system consisting of unidirectional transitions and demonstrates its tightness.
@@ VI Conclusions
We investigated the speed limits in discrete-time Markov chains. We found that time-reversed EP satisfies the speed limit in the same form as the conventional relation, while time-backward EP violates this bound. This contrasts with TUR: in a discrete-time process, the bound of time-reversed EP in TUR is exponentially suppressed or dependent on the minimum staying probability. Therefore, we anticipate that the discrete-time speed limit can provide a more efficient method for estimating EP in Markov chains compared to discrete-time TUR.
We also derived several variants of the conventional speed limit. The first is the stepwise speed limit which offers a bound in terms of distances between two consecutive probability distributions at all time steps. Second, we derived the section speed limit which is expressed solely by a single distance between any intermediate-time points. This relation is useful especially when only partial-time information on probability distributions is available. In addition, this relation can provide a nontrivial bound even in systems driven by a cyclic protocol and generally provides a tighter bound. Finally, we obtained a modified speed limit for systems with unidirectional transitions. This relation can be applied to various systems with absorbing states such as extinction dynamics and opinion dynamics.
@@ Acknowledgements.
This research was supported by the KIAS Individual Grant Nos. PG081802 (S.L.), and PG064902 (J.S.L.) at the Korea Institute for Advanced Study and an appointment to the JRG Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government (J.-M.P.). This was also supported by the Korean Local Governments - Gyeongsangbuk-do Province and Pohang City (J.-M.P.)

### v2

@@ V.3 Three-level system with unidirectional transition
$\mathbf{P}(t_{i+1})=\left(\mathbf{M}^{({\rm b})}+\mathbf{M}^{({\rm u})}\right)\cdot\mathbf{P}(t_{i}),$ (44)
@@
We evaluate time-reversed EP and dynamical activity for the one-time-step relaxation process described by Eq. (44). The results are plotted in Fig. 4. The dashed line represents the bound of the speed limit for a system including unidirectional transitions, as given by Eq. (40). Data points indicate the results obtained from various values of $g$, $M_{1,0}$, and $M_{0,1}$ chosen within the range $(0.01,0.9)$ under the constraint $d_{\textrm{TV}}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{(u)}(t_{0})>0$. This example validates the speed limit for a system consisting of unidirectional transitions and demonstrates its tightness.
V.4 Cell-state dynamics and theoretical computation
Finally, we consider two practical examples: cell-state dynamics Gupta et al. (2011) and a deterministic finite automaton Manzano et al. (2024). In Ref. Gupta et al. (2011), the dynamics of phenotypic states in human breast cancer cells were investigated experimentally. The study focused on three distinct differentiation states of cancer cells: stem-like, basal, and luminal. Observations revealed that the distribution of these cell states evolved over time during culture growth. To explain the observed dynamics, a discrete-time Markov chain model was employed.
We apply our results to the model for a human breast cancer cell line, known as SUM159, in Ref Gupta et al. (2011), where the transition matrix is given by
$\mathbf{M}^{\rm(159)}=\begin{pmatrix}0.58&0.01&0.04\\ 0.35&0.99&0.49\\ 0.07&0.00&0.47\\ \end{pmatrix}.$ (47)
Here the state distribution $P_{n}(t_{i})$ for $n=0,1,2$ represents the proportions of stem-like, basal, and luminal states, receptively, after $i$ days of growth in culture. The matrix for SUM 159 includes one unidirectional transition from the luminal state to the basal state. In Fig. 5, we plot the EP and the activity of bidirectional transitions for the one-step process of the breast cancer dynamics. The red ‘+’ symbols represent the data. The results are well bounded by our inequality.
In addition, we validate our speed limits on a finite automaton, a type of computational machine, designed to identify multiples of 4 introduced in Ref. Manzano et al. (2024). The computational machine has four states, $n=0,1,2,3$, and initially starts from $n=0$. Then, a given integer under identification is written in a string of binary numbers, either 00 or $1$, which is taken sequentially as an input. Depending on its current state and the input bit, the machine deterministically changes its state, which is designed to ensure that the final state is 00 if and only if the given integer is a multiple of 4. To investigate thermodynamics in this deterministic finite automaton, the authors considered a random input composed of independent and identically distributed random bits taking 00 with probability $p_{0}$ and $1$ with probability $p_{1}$. Then, the transition probability of the automaton is given by
$\mathbf{M}^{\rm(aut)}=\begin{pmatrix}p_{0}&0&p_{0}&0\\ p_{1}&0&p_{1}&0\\ 0&p_{0}&0&p_{0}\\ 0&p_{1}&0&p_{1}\end{pmatrix}.$ (48)
In Fig. 5, black circle symbols depict the numerical results of the automaton calculated in the one-step process with random initial conditions. Compared to the cell-state dynamics, the numerical data from the automaton dynamics spans a wider range of the plot.
Figure 5: Plot of the time-reversed EP versus the dynamical activity of bidirectional transitions for the one-step process of the breast cancer cell dynamics and the automaton dynamics in Sec. V.4. Both the $x$ and $y$ axes are scaled by the total variation distance minus unidirectional activity, $d_{\rm TV}[\mathbf{P}(t_{1}),\mathbf{P}(t_{0})]-A^{({\rm u})}(t_{0})$, for each point. The dashed line represents the speed limit for a system with unidirectional transitions, given by Eq. (40). The black circles indicate the calculation results obtained from the automaton’s dynamics by varying the initial condition, and the probabilities of its input, $p_{0}$ and $p_{1}$. The red ‘+’ symbols represent the results from the breast cancer cell dynamics of SUM 159 under various initial conditions with the transition matrix given by Eq. (47).
@@ VI Conclusions
We investigated the speed limits in discrete-time Markov chains. We found that the time-reversed EP satisfies the speed limit in the same form as the conventional relation, while time-backward EP violates this bound. These findings highlight significant differences from previous extensions of TUR to discrete-time processes. First, the conventional TUR is not valid for discrete-time processes, necessitating modifications such as the exponential TUR and the linear TUR Proesmans and Van den Broeck (2017); Liu et al. (2020). Second, both discrete-time TURs are not applicable to systems with time-dependent driving. In particular, the exponential TUR holds only in the steady state and the linear TUR holds for relaxation processes without time-dependent driving. Lastly, the bound of the exponential TUR rapidly decays as the system is driven further from equilibrium, and the linear TUR bound vanishes if at least one staying probability at any state is zero. In contrast, the examples in Sec. V demonstrate that our results hold for general time-dependent processes even in the presence of zero staying probabilities, providing a meaningful bound. Therefore, we anticipate that the discrete-time speed limit can offer a more efficient method for estimating EP in Markov chains compared to discrete-time TUR.
We also derived several variants of the conventional speed limit. The first is the stepwise speed limit which offers a bound in terms of distances between two consecutive probability distributions at all time steps. Second, we derived the interval speed limit which is expressed solely by a single distance between any intermediate-time points. This relation is useful especially when only partial-time information on probability distributions is available. In addition, this relation can provide a nontrivial bound even in systems driven by a cyclic protocol and generally provides a tighter bound. Finally, we obtained a modified speed limit for systems with unidirectional transitions. This relation can be applied to various systems with absorbing states such as extinction dynamics and opinion dynamics.
@@ Acknowledgements.
This research was supported by the KIAS Individual Grant Nos. PG081802 (S.L.), and PG064902 (J.S.L.) at the Korea Institute for Advanced Study and an appointment to the JRG Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government (J.-M.P.). This was also supported by the Korean Local Governments - Gyeongsangbuk-do Province and Pohang City (J.-M.P.). This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00557038) (J.-M. P.). S.L. acknowledges additional support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within SFB 1551 - project number 464588647. S.L. thanks Seung Ki Baek for valuable comments and references. S.L. also acknowledge the warm hospitality of Christopher Jarzynski at the Institute for Physical Science and Technology Department at the University of Maryland, College Park.
Appendix A Differences between $\Delta\Sigma$ and $\Delta\tilde{\Sigma}$
In this section, we compare $\Delta\Sigma$ and $\Delta\tilde{\Sigma}$ to describe their differences in several aspects.
A.1 Contributions of staying events
The transitions staying in the same state, occurring with probability $M_{n,n}(t_{i})$, gives no contributions to the time-reversed EP $\Delta\tilde{\Sigma}$ as in the continuous-time EP rate. In fact, $\Delta\tilde{\Sigma}$ have the same mathematical form as the continuous-time EP rate, given by
In contrast, the staying events contribute to time-backward EP.
A.2 Decomposition with system entropy change
The system entropy, defined as the Shannon entropy, is given by
$S_{\rm sys}(t_{i})\equiv-\sum_{n}P_{n}(t_{i})\ln P_{n}(t_{i}).$ (50)
We can decompose the time-backward EP into
$\Delta\Sigma(t_{i})=\Delta S_{\rm sys}(t_{i})+\Delta S_{\textrm{b}}(t_{i}),$ (51)
where $\Delta S_{\textrm{sys}}=S_{\textrm{sys}}(t_{i+1})-S_{\textrm{sys}}(t_{i})$ and
$\Delta S_{\textrm{b}}(t_{i})\equiv\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})}{M_{m,n}(t_{i})}.$ (52)
This decomposition is reminiscent of the thermodynamic EP rate for continuous-time processes, which is given by
$\dot{\Sigma}_{\textrm{cont}}(t)=$ $\dot{S}_{\text{sys}}(t)+\dot{S}_{\textrm{env}}(t),$ (53)
where the environmental entropy production rate is defined as
$\dot{S}_{\textrm{env}}(t)=\sum_{n,m}R_{n,m}(t)P_{m}(t)\ln\frac{R_{n,m}(t)}{R_{m,n}(t)}.$ (54)
For this reason, $\Delta S_{\textrm{b}}$ is referred to as the EP of the heat reservoirs in discrete-time process in Ref. Liu et al. (2020).
For the time-backward EP, the decomposition is given by
$\Delta\tilde{\Sigma}(t_{i})=\Delta S_{\rm sys}+\Delta\tilde{S}_{\textrm{b}}.$ (55)
with the entropy flow Gaspard (2004)
$\Delta\tilde{S}_{\textrm{b}}(t_{i})\equiv\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})P_{n}(t_{i+1})}{M_{m,n}(t_{i})P_{n}(t_{i})}.$ (56)
Unlike $\Delta S_{\textrm{b}}$ and the environmental EP in continuous-time processes, the summand in $\Delta\tilde{S}_{\textrm{b}}$ depends on $P_{n}(t)$.
A.3 Relation with detailed balance condition
Since time-reversed EP takes the form of Kullback-Leibler divergence, it vanishes if and only if $M_{n,m}(t_{i})P_{m}(t_{i})=M_{m,n}(t_{i})P_{n}(t_{i})$, which is called detailed balance condition in Ref. Ge et al. (2006). Since detailed balance ensures that the probability distribution remains unchanged, $P_{n}(t_{i+1})=P_{n}(t_{i})$, Eq. (6) shows that the time-backward EP also vanishes when the detailed balance is satisfied.
However, the detailed balance is not a necessary condition for vanishing time-backward EP. One simple example is a single bit under the NOT gate operation. Here, the bit system has two states $n=0,1$ and the transition probability matrix is given by
$\mathbf{M}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$ (57)
When the system starts from the initial condition $P_{n}(t_{0})=\delta_{n,0}$, the time-backward EP takes zero although the detailed balance condition is not obeyed in this process. However, the time-reversed EP is non-zero and even diverges. This example clearly demonstrates that the time-backward EP is not a proper indicator of the detailed balance condition.
The explicit condition for vanishing time-backward EP is $M_{n,m}(t_{i})P_{m}(t_{i})=M_{m,n}(t_{i})P_{n}(t_{i+1})$. By applying Bayes’ theorem, this condition can be rewritten as $M_{m,n}(t_{i})=M_{n,m}(t_{i})P_{m}(t_{i})/P_{n}(t_{i+1})=P(m,t_{i}|n,t_{i+1})$ with the conditional probability $P(m,t_{i}|n,t_{i+1})$ that the system was in the state $m$ at time $t_{i}$ given that it is in $n$ at time $t_{i+1}$. This condition reflects that the backward time evolution matches the forward evolution, $P(n,t_{i+1}|m,t_{i})=P(m,t_{i}|n,t_{i+1})$, and it can hold even when detailed balance is violated as shown in the NOT-gate example.
A.4 Relation with the H-theorem
In relaxation processes, where the transition probabilities are time-independent, the time-backward EP relates to the $H$-function, a monotonically decreasing quantity in time, defined as Morimoto (1963)
$H(t_{i})\equiv\sum_{n}P_{n}(t_{i})\ln\frac{P_{n}(t_{i})}{P_{n}^{\rm eq}}$ (58)
where $P_{n}^{\rm eq}$ is the equilibrium distribution. As shown in Lee (2018), the time-backward EP can be expressed as
$\Delta\Sigma(t_{i})=H(t_{i})-H(t_{i+1})\geq 0,$ (59)
indicating the connection between the time-backward EP and the $H$-theorem.
We note that it is a special case of the intrinsic mismatch cost, introduced in Ref. Kolchinsky and Wolpert (2017); Manzano et al. (2024), defined as
$\Delta\Sigma_{\textrm{miss}}(t_{i})=\sum_{n}P_{n}(t_{i})\ln\frac{P_{n}(t_{i})}{r_{n}(t_{i})}-\sum_{n}P_{n}(t_{i+1})\ln\frac{P_{n}(t_{i+1})}{r_{n}(t_{i+1})}$ (60)
with an arbitrary reference probability $r_{n}(t_{i})$ and $r_{n}(t_{i+1})=\sum_{m}M_{n,m}(t_{i})r_{n}(t_{i})$. If we choose $r_{n}(t_{i})=P_{n}^{\textrm{eq}}$, then it reduce to the time-backward EP.
Appendix B Derivation of Eq. (10)
In this section, we revisit Ref. Falasco et al. (2022) to represent the details for the derivation of the inequality in (10). To do this, we first derive two auxiliary relations. The first relation is an ordering between the total variation distance $d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]$ in Eq. (9) and the Le Cam’s distance defined as
$d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}]\equiv\sqrt{\frac{1}{2}\sum_{n}\frac{(p_{n}-q_{n})^{2}}{p_{n}+q_{n}}}.$ (61)
By using the Cauchy-Schwarz inequality, we can show that
$\left(2d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\right)^{2}$ $=\left(\sum_{n}|p_{n}-q_{n}|\right)^{2}$
$=\left(\sum_{n}\frac{|p_{n}-q_{n}|}{\sqrt{p_{n}+q_{n}}}\sqrt{p_{n}+q_{n}}\right)^{2}$
$\leq\left(\sum_{n}\frac{(p_{n}-q_{n})^{2}}{p_{n}+q_{n}}\right)\left(\sum_{n}(p_{n}+q_{n})\right)$
$=\left(2d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}]\right)^{2}.$ (62)
Since both distances are positive, it yields $d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\leq d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}]$.
To find the second relation, we express the squared Le Cam’s distance as
$\left(d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}]\right)^{2}=\frac{1}{2}\sum_{n}|p_{n}-q_{n}|\tanh\left(\frac{1}{2}\left|\ln\frac{p_{n}}{q_{n}}\right|\right).$ (63)
Here we used the property that $p_{n}-q_{n}$ and $\ln(p_{n}/q_{n})$ have the same sign. Since $\tanh x$ is concave for $x\geq 0$, by introducing a probability $r_{n}\equiv|p_{n}-q_{n}|/(2d_{\textrm{TV}}[\mathbf{p},\mathbf{q}])\geq 0$ and applying the Jensen’s inequality, we find
$\left(d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}]\right)^{2}$ $=d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\sum_{n}r_{n}\tanh\left(\frac{1}{2}\left|\ln\frac{p_{n}}{q_{n}}\right|\right)$
$\leq d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\tanh\left(\frac{1}{2}\sum_{n}r_{n}\left|\ln\frac{p_{n}}{q_{n}}\right|\right)$
$=d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\tanh\left(\frac{\sum_{n}(p_{n}-q_{n})\ln\frac{p_{n}}{q_{n}}}{4d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]}\right)$
$=d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\tanh\left(\frac{D[\mathbf{p}\parallel\mathbf{q}]+D[\mathbf{q}\parallel\mathbf{p}]}{4d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]}\right).$ (64)
Consequently, we arrive at
$(d_{\textrm{TV}}[\mathbf{p},\mathbf{q}])^{2}$ $\leq(d_{\textrm{LeCam}}[\mathbf{p},\mathbf{q}])^{2}$
$\leq d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\tanh\left(\frac{D_{\rm J}[\mathbf{p},\mathbf{q}]}{2d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]}\right).$ (65)
As $\tanh^{-1}x=(1/2)\ln((1+x)/(1-x))$, by rearranging this inequality, we finally obtain
$d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\ln\frac{1+d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]}{1-d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]}\leq D_{\rm J}[\mathbf{p},\mathbf{q}],$ (66)
which is equivalent to Eq. (10).